# Memory-Efficient Sequential Pattern Mining with Hybrid Tries

### v2

@@ 1. Phase 1:
    By the statement, all nodes $n^{\prime\prime}\in{\mathbf{P}}:n^{\prime\prime}\neq n^{\prime}$ have $\mathbb{I}_{n^{\prime\prime}}=\mathbb{I}_{n}$ and thus belong to the same itemset. By definition, events within an itemset are unique and cannot be repeated. Therefore, we have $\mathbb{I}_{a_{n^{\prime\prime}}}<\mathbb{I}_{n}$, which by Theorem 11, means that all nodes $n^{\prime\prime}$ model an itemset extension $\langle\mathcal{P},e_{n^{\prime\prime}}$.
    Similarly, as node $n^{\prime}$ is the first node in ${\mathbf{P}}$ such that $\mathbb{I}_{n^{\prime}}\neq\mathbb{I}_{n}$, its ancestor $a_{n^{\prime}}$ must either occur prior to node $n$ or within the same itemset as node $n$. This means $\mathbb{I}_{a_{n^{\prime}}}\leq\mathbb{I}_{n}$, which by Theorem 11, means that node $n^{\prime}$ models a sequence extension $\langle\mathcal{P},e_{n^{\prime}}$. ∎
@@ 2. Phase 2:
In the first iteration, $\mathcal{BT}$Miner initializes with pattern-node pairs $(\langle\bar{a}\rangle,\{n_{1}\}),(\langle\bar{b}\rangle,\{n_{2},n_{7},n_{16}\}),(\langle\bar{c}\rangle,\{n_{5},n_{8},n_{11},n_{15}\})$. The mining algorithm takes a pattern-node pair, for example, $(\langle\bar{a}\rangle,\{n_{1}\})$, and in the first phase performs a depth-first-search of the subtrie rooted at node $n_{1}$ to find the first nodes $n^{\prime}:\mathbb{I}_{n^{\prime}}\neq\mathbb{I}_{n}$. This results in finding nodes $n_{3},n_{6},n_{7},n_{9}$. All these node model a sequence extension according to Lemma 12, and give a pattern-node pair $\left(\langle\bar{a},\bar{a}\rangle,\{n_{3},n_{9}\}),(\langle\bar{a},\bar{b}\rangle,\{n_{7}\}),(\langle\bar{a},\bar{c}\rangle,\{n_{6}\})\right)$. On the other hand, any node traversed on the paths ${\mathbf{P}}=(n_{1},\dots,n^{\prime}),n^{\prime}\in\{n_{3},n_{6},n_{7},n_{9}\}$ models an itemset extension according to Lemma 12, which gives pattern-node pairs $(\langle\bar{a},b\rangle,\{n_{2}\})(\langle\bar{a},c\rangle,\{n_{5},n_{8}\})$.
@@
The worst-case space complexity of $\mathcal{BT}$Miner is given by Theorem 15
@@
As proved in Lemma 5, trie models of the data set are identical in their node set $\mathcal{N}$, and by Lemma 6, the number of nodes in a trie model is $\mathcal{O}\left(\min\{N,|\mathbb{E}|^{M}\}M\right)$. This gives the total worst-case memory complexity of $\mathcal{BT}$ as $\mathcal{O}\left(\min\{N,|\mathbb{E}|^{M}\}M\log(N)\right)$.
@@ 3.3 Hybrid Tries and the $\mathcal{HT}$Miner Algorithm
Regardless of the overlap for sequences $\mathcal{S}_{i}\in\mathcal{SD}$, sequential data sets generally have a high overlap on their prefixes $\mathcal{\check{S}}_{i}:\mathcal{\check{S}}_{i}\sqsubseteq\mathcal{S}_{i}\in\mathcal{SD}$. This is due to the lower number of possible event combinations $\mathbb{E}^{j}\geq|\mathcal{N}_{j}|$ at any length $j\leq|\mathcal{S}_{i}|$, which is typically much lower than $N$ for smaller values of $j$. For such prefixes, a trie model is more memory efficient. As $j$ increases, so does the possible number of unique and non-overlapping sequences, leading to potentially less overlap on postfixes $\mathcal{\hat{S}}_{i}\sqsubseteq\mathcal{S}_{i}\in\mathcal{SD}$. For such postfixes, a vector model is more memory efficient in practice. We propose a hybrid model $\mathcal{HT}$ to take advantage of this trade-off.
@@
An important challenge of our hybrid model is determining the transition length $j^{*}$ that provides the highest memory efficiency. This is data set dependent and determined based on the trade-off between the compression that a $\mathcal{BT}$ model can provide for layers $\mathcal{N}_{j},0<j\leq j^{*}$, versus the memory overhead introduced by the transition to a vector model for postfixes $\mathcal{\hat{S}}_{i}:|\mathcal{\hat{S}}_{i}|=M-j^{*}$. In particular, we choose $j^{*}$ using Theorem 16. Here, $N_{j}$ denotes the number of sequences $\mathcal{S}_{i}\in\mathcal{SD}:|\mathcal{S}_{i}|\geq j$, recorded during the preprocessing step of constructing $\mathcal{BT}$.
@@
For an $\mathcal{HT}$ transitioning at a layer $\mathcal{N}_{j^{\prime}}\in\mathcal{HT}:j^{\prime}\in\{0,\dots,M$ involves a $\mathcal{BT}$ model of all prefixes $\mathcal{\check{S}}_{i}\sqsubseteq\mathcal{S}_{i}\in\mathcal{SD}:|\mathcal{\check{S}}_{i}|=j^{\prime}$, and a vector model of postfixes $\mathcal{\hat{S}}_{i}\sqsubset\mathcal{S}_{i}:|\mathcal{\hat{S}}_{i}|=|\mathcal{S}_{i}|-j^{\prime}$.
@@
[…]
@@ 7. $j=$ 6:
4:     Mine the vector $\mathcal{V}$ pointed to by $i$ starting from position $j$ according to prefix projection.
@@
10:              Mine the vector $\mathcal{V}$ starting from position $1$ according to prefix projection.
@@
13:              Mine the vector $\mathcal{V}$ starting from position $1$ according to prefix projection.
@@
To mine $\mathcal{HT}$ models, we integrate $\mathcal{BT}$Miner and prefix projection into the $\mathcal{HT}$Miner algorithm. The $\mathcal{HT}$Miner algorithm mines its trie model of prefixes $\mathcal{\check{S}}_{i}$ using the procedure of $\mathcal{BT}$Miner, and its vector model of postfixes $\mathcal{\hat{S}}_{i}$ using prefix projection. If the mining algorithm traverses a transitioning node $n\in\mathcal{N}_{j^{*}}$, it uses the ancestor vector $\vec{a}$ and Theorem 11 to determine valid pattern extensions in the vector model. The complete procedure is given in Algorithm 4.
@@ 4 Numerical Results
For our numerical tests, we evaluated the effect of different data set models on the performance of state-of-the-art mining algorithms in large-scale SPM. We use PrefixSpan by Han et al. (2001), which forms the basis of almost all state-of-the-art vector-based prefix-projection algorithms, and TreeMiner (Rizvee et al., 2020), which is the state-of-the-art trie model for SPM. Note that TreeMiner uses several additional mining techniques, such as co-occurrence information of events (Fournier-Viger et al., 2014) for faster SPM, which are not implemented in our mining algorithms. Such algorithmic enhancements, or any other enhancements for time efficiency in the rich literature of SPM can also be implemented on our models without loss of generality.
All algorithms were coded in C++ and executed on a PC with an Intel Xeon W-2255 processor, 256GB of memory, and Ubuntu 20.04.1 operating system. [^1] Our algorithms are open source and available at https://github.com/aminhn/BDTrie Note that our system memory is considerably higher than an average PC, often capped at 32GB. We limit all tests to one core of the CPU and a 36000-second time limit. All instances are mined with minimum support thresholds of $\theta\in\{10\%,8\%,6\%,4\%,2\%,0.5\%\}$.
@@ 4.2 Memory Consumption
Figure 5 gives the peak memory consumption of all algorithms. In the small data sets of Kosarak and MSNBC, all algorithms consume less than 1GB of memory, with the exception of TreeMiner in the Kosarak data set, which uses close to 60GB of memory. This is due to the higher number of events $|\mathbb{E}|$ in Kosarak, which considerably affects the memory usage of TreeMiner due to its set of labels $\mathcal{L}$. In the Kosarak data set, $\mathcal{BT}$Miner and $\mathcal{HT}$Miner use less than 0.1GB of memory, an average improvement of 86% over PrefixSpan and 99.9% over TreeMiner. In the MSNBC data set, $\mathcal{BT}$Miner and $\mathcal{HT}$Miner show an average improvement of 66% over PrefixSpan and 94% over TreeMiner. $\mathcal{HT}$Miner is mostly similar to $\mathcal{BT}$Miner on both data sets, but shows an average improvement of 30% in the MSNBC data set.
@@
In the medium data set Twitch, PrefixSpan uses less than 9GB of memory, $\mathcal{BT}$Miner uses less than 7GB of memory with an average improvement of 79%, and $\mathcal{HT}$Miner uses less than 2GB of memory with an average improvement of 91%. The improvement of $\mathcal{HT}$Miner is higher for lower support thresholds. TreeMiner exceeds the 256GB system memory and can only fit into memory 0.6% of the Twitch data set. In the Spotify data set, the trie models show more than two orders of magnitude memory saving over the vector model of PrefixSpan. In particular, PrefixSpan uses 10GB of memory while TreeMiner uses 0.12GB of memory with an average improvement of 98.8%. The $\mathcal{BT}$Miner and $\mathcal{HT}$Miner algorithms use less than 0.02GB of memory and show improvements of 99.7% over PrefixSpan and 78% over TreeMiner. The $\mathcal{BT}$Miner and $\mathcal{HT}$Miner are equivalent in the Spotify data set as the hybrid algorithm determined the transition length as $j^{*}=M$, giving a full trie model of the data set for $\mathcal{HT}$Miner.
@@
The $\mathcal{BT}$Miner and $\mathcal{HT}$Miner algorithms enjoy the most significant improvements in large data sets. In the Criteo data set, PrefixSpan exceeds the 256GB of system memory and is only able to model 15% of the data set. Although TreeMiner can model the entire Criteo data set, it exceeds system memory during its mining algorithm and terminates. This is in contrast to $\mathcal{BT}$Miner and $\mathcal{HT}$Miner, which use at least 2GB and at most 25GB of memory to mine the entire data set. This is more than an order of magnitude memory saving, and amounts to at least 1,700GB lower memory usage than PrefixSpan requires to model the full data set into memory. Similar to the Spotify data set, $\mathcal{BT}$Miner and $\mathcal{HT}$Miner are equivalent with $j^{*}=M$.
@@
Overall, $\mathcal{HT}$Miner is the most memory-efficient SPM algorithm, closely followed by $\mathcal{BT}$Miner. TreeMiner is the least memory-efficient algorithm, using considerably higher memory than all other algorithms and often failing to model medium to large data sets within system memory. While less efficient, the memory usage of PrefixSpan is still within the range of what is typical for regular PCs for small to medium data sets. However, this is not the case for large data sets, with $\mathcal{HT}$Miner being the only capable SPM algorithm. We conclude that $\mathcal{HT}$Miner provides compelling benefits for any data set size over all other alternatives.

### v3

@@ 1. Phase 1:
    By the statement, all nodes $n^{\prime\prime}\in{\mathbf{P}}:n^{\prime\prime}\neq n^{\prime}$ have $\mathbb{I}_{n^{\prime\prime}}=\mathbb{I}_{n}$ and thus belong to the same itemset. By definition, events within an itemset are unique and cannot be repeated. Therefore, we have $\mathbb{I}_{a_{n^{\prime\prime}}}<\mathbb{I}_{n}$, which by Theorem 10, means that all nodes $n^{\prime\prime}$ model an itemset extension $\langle\mathcal{P},e_{n^{\prime\prime}}\rangle$.
    Similarly, as node $n^{\prime}$ is the first node in ${\mathbf{P}}$ such that $\mathbb{I}_{n^{\prime}}\neq\mathbb{I}_{n}$, its ancestor $a_{n^{\prime}}$ must either occur prior to node $n$ or within the same itemset as node $n$. This means $\mathbb{I}_{a_{n^{\prime}}}\leq\mathbb{I}_{n}$, which by Theorem 10, means that node $n^{\prime}$ models a sequence extension $\langle\mathcal{P},e_{n^{\prime}}\rangle$. ∎
@@ 2. Phase 2:
In the first iteration, $\mathcal{BT}$Miner initializes with pattern-node pairs $(\langle\bar{a}\rangle,\{n_{1}\}),(\langle\bar{b}\rangle,\{n_{2},n_{7},n_{16}\}),(\langle\bar{c}\rangle,\{n_{5},n_{8},n_{11},n_{15}\})$. The mining algorithm takes a pattern-node pair, for example, $(\langle\bar{a}\rangle,\{n_{1}\})$, and in the first phase performs a depth-first-search of the subtrie rooted at node $n_{1}$ to find the first nodes $n^{\prime}:\mathbb{I}_{n^{\prime}}\neq\mathbb{I}_{n}$. This results in finding nodes $n_{3},n_{6},n_{7},n_{9}$. All these nodes model a sequence extension according to Lemma 11, and give a pattern-node pair $\left(\langle\bar{a},\bar{a}\rangle,\{n_{3},n_{9}\}),(\langle\bar{a},\bar{b}\rangle,\{n_{7}\}),(\langle\bar{a},\bar{c}\rangle,\{n_{6}\})\right)$. On the other hand, any node traversed on the paths ${\mathbf{P}}=(n_{1},\dots,n^{\prime}),n^{\prime}\in\{n_{3},n_{6},n_{7},n_{9}\}$ models an itemset extension according to Lemma 11, which gives pattern-node pairs $(\langle\bar{a},b\rangle,\{n_{2}\})(\langle\bar{a},c\rangle,\{n_{5},n_{8}\})$.
@@
The worst-case space complexity of $\mathcal{BT}$Miner is given by Theorem 14.
@@
As trie models of the data set are identical in their node set $\mathcal{N}$, and by Lemma 5, the number of nodes in a trie model is $\mathcal{O}\left(\min\{N,|\mathbb{E}|^{M}\}M\right)$. This gives the total worst-case memory complexity of $\mathcal{BT}$ as $\mathcal{O}\left(\min\{N,|\mathbb{E}|^{M}\}M\log(N)\right)$.
@@ 3.3 Hybrid Tries and the $\mathcal{HT}$Miner Algorithm
Regardless of the overlap for sequences $\mathcal{S}_{i}\in\mathcal{SD}$, sequential data sets generally have a high overlap on their prefixes $\mathcal{\check{S}}_{i}:\mathcal{\check{S}}_{i}\sqsubseteq\mathcal{S}_{i}\in\mathcal{SD}$. This is due to the lower number of possible event combinations $|\mathbb{E}|^{j}\geq|\mathcal{N}_{j}|$ at any length $j\leq|\mathcal{S}_{i}|$, which is typically much lower than $N$ for smaller values of $j$. For such prefixes, a trie model is more memory efficient. As $j$ increases, so does the possible number of unique and non-overlapping sequences, leading to potentially less overlap on postfixes $\mathcal{\hat{S}}_{i}\sqsubseteq\mathcal{S}_{i}\in\mathcal{SD}$. For such postfixes, a vector model is more memory efficient in practice. We propose a hybrid model $\mathcal{HT}$ to take advantage of this trade-off.
@@
An important challenge of our hybrid model is determining the transition length $j^{*}$ that provides the highest memory efficiency. This is data set dependent and determined based on the trade-off between the compression that a $\mathcal{BT}$ model can provide for layers $\mathcal{N}_{j},0<j\leq j^{*}$, versus the memory overhead introduced by the transition to a vector model for postfixes $\mathcal{\hat{S}}_{i}:|\mathcal{\hat{S}}_{i}|=M-j^{*}$. We choose $j^{*}$ using Theorem 15, where $N_{j}$ denotes the number of sequences $\mathcal{S}_{i}\in\mathcal{SD}:|\mathcal{S}_{i}|\geq j$, recorded during the preprocessing step of constructing $\mathcal{BT}$.
@@
For an $\mathcal{HT}$ transitioning at a layer $\mathcal{N}_{j^{\prime}}\in\mathcal{HT}:j^{\prime}\in\{0,\dots,M\}$ involves a $\mathcal{BT}$ model of all prefixes $\mathcal{\check{S}}_{i}\sqsubseteq\mathcal{S}_{i}\in\mathcal{SD}:|\mathcal{\check{S}}_{i}|=j^{\prime}$, and a vector model of postfixes $\mathcal{\hat{S}}_{i}\sqsubset\mathcal{S}_{i}:|\mathcal{\hat{S}}_{i}|=|\mathcal{S}_{i}|-j^{\prime}$.
@@
We have $N_{j}=\langle 20,19,16,10,4,3\rangle$, and $|\mathbb{E}|_{j}=\langle 1,3,2,3,3,2\rangle$. For our system, we have $c^{\text{int}}=4,c^{\mathcal{V}}=24$, and by Theorem 14, $c^{\mathcal{T}}=6$. Checking the equation of Theorem 15, we have the following memory estimates for $j\in\{0,\dots,6\}$:
@@ 7. $j=$ 6:
4:     Mine the vector $\mathcal{V}$ pointed to by $i$ starting from position $j$ according to prefix-projection.
@@
10:              Mine the vector $\mathcal{V}$ starting from position $1$ according to prefix-projection.
@@
13:              Mine the vector $\mathcal{V}$ starting from position $1$ according to prefix-projection.
@@
To mine $\mathcal{HT}$ models, we combine $\mathcal{BT}$Miner and prefix-projection into a novel $\mathcal{HT}$Miner algorithm. The $\mathcal{HT}$Miner algorithm mines its trie model of prefixes $\mathcal{\check{S}}_{i}$ using the procedure of $\mathcal{BT}$Miner, and its vector model of postfixes $\mathcal{\hat{S}}_{i}$ using prefix-projection. If the mining algorithm traverses a transitioning node $n\in\mathcal{N}_{j^{*}}$, it uses the ancestor vector $\vec{a}$ and Theorem 10 to determine valid pattern extensions in the vector model. The complete procedure is given in Algorithm 4.
@@ 4 Numerical Results
For our numerical tests, we evaluated the effect of different data set models on the performance of state-of-the-art mining algorithms in large-scale SPM. We use PrefixSpan by Han et al. (2001), which forms the basis of almost all state-of-the-art vector-based prefix-projection algorithms, and TreeMiner (Rizvee et al., 2020), which is the only available trie model for SPM. Note that TreeMiner uses several additional mining techniques, such as co-occurrence information of events (Fournier-Viger et al., 2014) for faster SPM, which are not implemented in our mining algorithms for a base-case comparison. Nonetheless, any such algorithmic enhancement developed for time efficiency in the rich literature of SPM can also be implemented on our models without loss of generality.
All algorithms were coded in C++ and executed on a PC with an Intel Xeon W-2255 processor, 256GB of memory, and Ubuntu 20.04.1 operating system. [^1] Our algorithms are open source and available at https://github.com/aminhn/BDTrie Note that our system memory is considerably higher than an average PC, often ranging between 8-32GB. We limit all tests to one core of the CPU and a 36,000-second time limit.
@@ 4.1 Data Sets
Table 3: Number of frequent patterns by data set and support threshold.
Criteo (large)
$\theta$ 10% 8% 6% 4% 2% 0.5%
Patterns found 10,705 20,132 (41,559) (78,598) (88,434) (182,541)
Genome (large)
$\theta$ 10% 8% 6% 4% 2% 0.5%
Patterns found (264) (314) (459) (552) (598) (615)
Twitch (medium)
$\theta$ 8% 4% 2% 1% 0.5% 0.15%
Patterns found 2 14 160 1,382 13,493 789,685
Spotify (medium)
$\theta$ 1% 1e-1% 1e-2% 1e-3% 1e-4% 1e-5%
Patterns found 59 377 2,559 16,253 107,257 696,098
Kosarak (small)
$\theta$ 4% 2% 1% 0.5% 0.25% 0.1%
Patterns found 29 94 329 1,462 8,427 758,141
MSNBC (small)
$\theta$ 3% 1% 0.3% 0.1% 0.03% 0.01%
Patterns found 22 254 2,014 16,620 303,917 102,108,060
Note. Parenthesis represent lower bound.
@@
An important aspect of mining the above data sets is the imposed support threshold $\theta$. Higher support thresholds lead to fewer frequent patterns, mainly showcasing the time and memory efficiency of building and storing a data set representation in memory. Lower support thresholds lead to a larger number of frequent patterns, mainly showcasing the time and memory efficiency of a mining algorithm. We choose a wide range of thresholds for each data set tailored to observe and showcase both scenarios. The thresholds and number of frequent patterns found for each data set are given in Table 3. Note that when all algorithms exceeded the imposed time, the reported number of frequent patterns is a lower bound.
@@ 4.2 Memory Consumption
Figure 5 gives the peak memory consumption of all algorithms. In the small data set Kosarak, all algorithms, with the exception of TreeMiner, consume less than 1GB of memory. TreeMiner uses close to 57GB of memory, mainly due to the higher number of events $|\mathbb{E}|$ in this data set which is detrimental to its memory efficiency. The $\mathcal{HT}$Miner algorithm uses less than 0.7GB of memory which amounts to an average improvement of 68% (less than 0.2GB) over PrefixSpan and 99.9% (more than 56GB) over TreeMiner. In contrast to all other data sets, $\mathcal{HT}$Miner uses slightly higher memory (0.1GB on average, and 0.4GB at most) compared to $\mathcal{BT}$Miner in lower support thresholds. This is mainly due to the heuristic selection of the transitioning layer, where values $|\mathcal{N}_{j}|$ were overestimated, leading to a lower than optimal value $j^{*}$.
In the small data set MSNBC, $\mathcal{HT}$Miner shows an average improvement of 60% (less than 0.1GB) over PrefixSpan and 93% (approximately 0.5GB) over TreeMiner. Similarly, $\mathcal{BT}$Miner shows an average improvement of 37% (less than 0.1GB) over PrefixSpan and 89% (approximately 0.5GB) over TreeMiner. These improvements show the slight advantage of compactly fitting the data in memory using a trie with efficient labels, even for smaller data sets. $\mathcal{HT}$Miner shows an average improvement of 36% (less than 0.1GB) over $\mathcal{BT}$Miner, mainly due to the longer sequences in the MSNBC data set which are more efficiently modeled by a hybrid trie structure.
In the medium data set Twitch, PrefixSpan uses less than 16GB of memory, $\mathcal{BT}$Miner uses approximately 11GB of memory with an average improvement of 64% over PrefixSpan, and $\mathcal{HT}$Miner uses less than 3GB of memory with an average improvement of 89% over PrefixSpan and 54% over $\mathcal{BT}$Miner. The improvement of $\mathcal{HT}$Miner and $\mathcal{BT}$Miner is higher for larger support thresholds due to a more compact representation of the data set. TreeMiner exceeds the 256GB system memory and can only fit into memory only 0.6% of the Twitch data set. This is again mainly due to the high number of events $|\mathbb{E}|$ which negatively affects the memory efficiency of TreeMiner.
@@
In the Spotify data set, the trie models show more than two orders of magnitude memory saving over the vector model of PrefixSpan. In particular, PrefixSpan uses approximately 10GB of memory while TreeMiner uses 0.12GB of memory with an average improvement of 98.8%. The $\mathcal{BT}$Miner and $\mathcal{HT}$Miner algorithms use less than 0.02GB of memory and show improvements of 99.7% over PrefixSpan and 78% over TreeMiner. These results are mainly due to the low number of events $|\mathbb{E}|$ in the Spotify data set, which allows higher compression in trie models. The $\mathcal{BT}$Miner and $\mathcal{HT}$Miner are equivalent in the Spotify data set as the hybrid algorithm determined the transition length as $j^{*}=M$, giving a full trie model of the data set for $\mathcal{HT}$Miner. Both algorithms are more efficient than TreeMiner due to their lower memory overhead at each trie node.
@@
The $\mathcal{BT}$Miner and $\mathcal{HT}$Miner algorithms enjoy the most significant improvements in large data sets. In the Criteo data set, PrefixSpan exceeds the 256GB of system memory and is only able to model 15% of the data set. Although TreeMiner can model the entire Criteo data set, it exceeds system memory during its mining algorithm and terminates. This is in contrast to $\mathcal{BT}$Miner and $\mathcal{HT}$Miner, which use at least 2GB and at most 25GB of memory to mine the entire data set. This is more than an order of magnitude memory saving, and potentially amounts to 1.7TB lower memory usage than PrefixSpan requires to model the full data set into memory. This showcases the strength of trie models in modeling overlapping sequences using only a single path, while vector models use multiple vectors. Similar to the Spotify data set, $\mathcal{BT}$Miner and $\mathcal{HT}$Miner are equivalent with $j^{*}=M$.
@@
Overall, $\mathcal{HT}$Miner is the most memory-efficient SPM algorithm, closely followed by $\mathcal{BT}$Miner. TreeMiner is the least memory-efficient algorithm, using considerably higher memory than all other algorithms and often failing to model medium to large data sets within system memory. While less efficient, the memory usage of PrefixSpan is still within the range of what is typical for regular PCs for small to medium data sets. However, this is not the case for large data sets, with $\mathcal{HT}$Miner being the only capable SPM algorithm. We conclude that $\mathcal{HT}$Miner provides compelling benefits in memory efficiency for any data set size over all other alternatives.